# SPIN-ODE: Stiff Physics-Informed Neural ODE for Chemical Reaction Rate Estimation

### v2

@@ III-B Proposed method
where $y_{\text{max}}$ and $y_{\text{min}}$ are the species-wise maximum and minimum concentrations used in the normalisation layer, and $t_{\text{scale}}$ is the time span of the entire measurements. The ODE solver integrates the output time derivative to get a trajectory as Eq (9), and the loss function is correspondingly normalised by the scale of concentrations as:
@@
In addition to the loss on predicted state variables, we add two physical loss terms on the velocity and acceleration of the concentration change, enabling the learned dynamics to be further constrained by the physical laws. Specifically, we compute the 1st-order and 2nd-order derivatives losses ($\mathcal{L}_{\dot{y}}$ and $\mathcal{L}_{\ddot{y}}$) with respect to time, then apply the same scaling technique as in Equation (11), and the final loss function is:
$\mathcal{L}=\mathcal{L}_{y}+\mathcal{L}_{\dot{y}}+\mathcal{L}_{\ddot{y}}.$ (12)
SPIN-ODE step 2: Resampling and pre-training CRNN. While non-linear system dynamics can often be accurately modelled using simple deep neural networks, i.e., MLPs, the parameters learned by such black-box models typically lack interpretability. In contrast, physics-based neural ODEs encode parameters and mathematical operators derived directly from physical equations [38, 43]. Following [38, 39], we leverage the Chemical Reaction Neural Network (CRNN) (see Figure 2), a physics-based neural network tailored for discovering chemical reaction parameters. CRNN expresses the reaction rate law in Eq (2) as a linear combination of the logarithm of the reaction rate coefficient and the logarithms of species concentrations, followed by an exponential transformation. Each network node represents an individual reaction, and nodes can be stacked to form a comprehensive model of a reaction system. The overall time derivative of species concentrations $d\mathbf{Y}/dt$, is obtained by linearly transforming the reaction rate vector $\mathbf{R}$, through the net stoichiometric matrix $\mathbf{S}$, defined as $s_{i,j}=-s_{i,j}^{f}+s_{i,j}^{r}$:
$r_{i}=\exp\left(\ln(k_{i})+s_{i,a}^{f}\ln[A]+s_{i,b}^{f}\ln[B]+s_{i,c}^{f}\ln[C]\right).$ (13)
@@
To facilitate CRNN training, we first perform dense sampling of data points from the pre-trained MLPs in step 1, and then use the resulting “augmented points” for supervised learning. Specifically, when the latent neural ODE fits the concentration trajectory ${{y}(t_{i})}_{i=1}^{M}$ with original $M$ points, we increase the time resolution and integrate an interpolated concentration trajectory ${\hat{y}(t_{i})}_{i=1}^{N}$ with $N>M$. We then apply the Central Difference Formula (CDF) to estimate the corresponding time derivative $\dot{\hat{y}}(t)$ at state $\hat{y}(t)$:
$\dot{\hat{y}}(t_{i})\approx\frac{\hat{y}(t_{i+1})-\hat{y}(t_{i-1})}{t_{i+1}-t_{i-1}},\quad\text{for }1<i<N,$ (15)
with forward and backward difference schemes applied at the endpoints $i=1$ and $i=N$ respectively:
$\dot{\hat{y}}(t_{1})$ $\approx\frac{\hat{y}(t_{2})-\hat{y}(t_{1})}{t_{2}-t_{1}},$ (16)
$\dot{\hat{y}}(t_{N})$ $\approx\frac{\hat{y}(t_{N})-\hat{y}(t_{N-1})}{t_{N}-t_{N-1}}.$ (17)
Using this augmented dataset of $\left(\hat{y}(t_{i}),\dot{\hat{y}}(t_{i})\right)$, we can formulate a supervised learning task using the same scaled MSE loss as Eq 11.
The reason behind such ”first-regression-then-estimation“ design is twofold: 1) the physical interpretability of network parameters depends on their alignment with the corresponding physical equations, thus precluding operations that violate physical laws; and 2) for non-linear chemical kinetics, we apply linear normalisation directly to input concentrations and denormalisation in step 1 for stable optimisation. It induces non-linear distortions of reaction rates, causing deviations of learned parameters from their true physical meaning.
SPIN-ODE step 3: Fine-tuning CRNN with ODE solver. After the CRNN is pre-trained, we resume the training by combining the CRNN and the ODE solver to fine-tune the rate coefficients. The loss function is as the same as that in Eq 12. The initial learned rate coefficients during step 2 avoid the extreme error growth and stabilise the ODE integration. We finally extract the learned rate coefficients $k_{i}$ in the physics-based CRNN as the final estimation.
IV Experiments
@@
SPIN-ODE-s1 $2.8\times 10^{-3}$ $1.1\times 10^{-5}$ $1.4\times 10^{-5}$
@@
TABLE I: Overall comparisons between our proposed three-step SPIN-ODE approach and Ji et al. [38] model. We show the scaled MSE loss of the ODE trajectory fitting, as well as the log-MAE loss on rate coefficient estimation. - means the network involved in the current step does not provide estimation, while * means runtime error.
IV-A Experimental setups
Our method is first validated using the Robertson problem, a classical benchmark comprising three chemical species and three highly stiff reactions as presented in Eq (1). The governing equations for the Robertson problem are explicitly defined in Eq (5), with ground-truth reaction rate coefficients as $[0.04,3\times 10^{7},1\times 10^{4}]$.
Real-world atmospheric chemistry systems typically involve significantly more complex dynamics, exhibiting substantial stiffness across numerous reactions. Consequently, we further evaluate our method on the POLLU problem [44], a widely utilised atmospheric chemistry model for stiff ODE benchmarks. This system includes 25 reactions among 20 chemical species, with reaction rate coefficients spanning from $10^{-3}$ to $10^{12}$. In addition, we introduce an autoxidation reaction (AOXID) scheme involving 44 species and 49 reactions. It is a simplified system representing realistic autoxidation processes in the atmosphere characterized by intricate radical chain mechanisms [45]. The reaction rate coefficients in this system vary dramatically, ranging from approximately $10^{-14}$ to $10^{6}$. Comprehensive details for the POLLU and AOXID reaction schemes are provided in the supplementary materials.
Datasets for the Robertson and POLLU problems are numerically generated using a stiff-aware Backward Differentiation Formula (BDF) solver available in the SciPy library. The AOXID dataset is generated using the Rosenbrock solver within the well-known Kinetic PreProcessor (KPP) [46]. For the Robertson system, initial concentrations are set as $[y_{1},y_{2},y_{3}]=[1,0,0]$, with 50 data points sampled logarithmically spaced between $10^{-5}$ and $10^{5}$ seconds. For the POLLU and AOXID datasets, initial conditions detailed in the supplementary material are used, and 100 data points are uniformly sampled across the time span of 0.1 and 100 seconds, respectively. To mitigate numerical instabilities arising from prolonged integration and accumulation of errors, a sliding window approach of size 20 and stride of 10 is adopted during the training for POLLU and AOXID.
The proposed SPIN-ODE leverages the stiff-aware differentiable ODE solver Kvaerno3 from the diffrax library [47]. The MLP module is based on a 3-layer Multilayer Perceptron with 128 nodes per hidden layer and GELU activation. Experiments are conducted using an NVIDIA A100 GPU with 40 GB of memory. Network parameters are optimised using the Adam algorithm, initialised with a learning rate of 0.001, which is annealed when training loss does not decrease for 10% of the total epochs. The code is available at https://github.com/pvvq/SPIN-ODE.
IV-B Comparison with state of the art
Our overall comparisons include both ODE trajectory fitting and rate coefficient estimation. Given the wide-ranging magnitudes of the chemical concentration and reaction rate coefficients, we evaluate the accuracy of model estimations using the mean absolute error (MAE) on a logarithmic scale. To the best of our knowledge, we are the first to study stiff chemical ODE for rate coefficient estimation. There is only one related work, Ji et al. [38], which we reimplement for comparisons.
The quantitative results on ODE fitting and rate coefficients are reported in Table I. We show our results in every step (SPIN-ODE-s1, SPIN-ODE-s2, SPIN-ODE-s3). As shown in Table I, ours outperforms Ji et al. [38] in every step. It is interesting to see that Ji et al. fail to converge on the AOXID dataset, and large error on the POLLU dataset. On the other hand, our final results (SPIN-ODE-s3) achieve low errors in rate coefficient estimation and ODE fitting, specifically $7.4\times 10^{-3}$ for Robertson, $8.6\times 10^{-6}$ for POLLU, and $4.5\times 10^{-6}$ for AOXID. The promising results on the AOXID dataset indicate that our approach could be used in real-world atmospheric chemistry. Furthermore, by comparing SPIN-ODE-s3 with SPIN-ODE-s2, it demonstrates the benefit from the final refinement process. Note that in SPIN-ODE-s1, we use MLPs to implicitly learn ODE fitting so that it cannot estimate rate coefficients yet. Comparing SPIN-ODE-s1 and SPIN-ODE-s2, the latter increases the ODE fitting errors because it learns to search for good rate coefficients to fit the physics-based CRNN model. The CRNN has fewer free parameters than MLPs to minimize the overall ODE fitting errors. However, once the optimal values of $k$ is found, SPIN-ODE-s3 produces the lowest errors in the ODE fitting.
To better visualise the ODE trajectory fitting, we plot the trajectory predictions in Figure 3. From left to right, we have the results for Robertson, POLLU and AOXID. Note that we select the three most representative reactions from POLLU and AOXID for visualization. The complete reaction plots can be found in the supplementary material. Since the method by Ji et al. performs reliably only on the Robertson problem, we did not show its trajectories on POLLY and AOXID. For our approach, it fits the ground truth trajectories in all three cases.
Since our primary focus is on rate coefficient estimation, we also plot the scatter diagram to demonstrate all rate coefficients in Figure 4. We can see that the model predictions (blue cross) closely match with ground truth values (yellow circles). Note that for the POLLU problem, coefficients 15-19 correspond to oxidant cycling, whose rate coefficients have been intensively studied and are provided during training and thus excluded from the estimation target.
@@ IV-B Comparison with state of the art
IV-C Ablation studies
@@
We compare four strategies for our pre-training step to illustrate the proposed key components: 1) we directly apply the Central Difference Formula (CDF) to the original observation and obtain the time derivative without SPIN-ODE-s1 MLP fitting, we refer to it as Direct CDF, 2) after MLP fitting, we directly use MLP outputs to train the CRNN, without SPIN-ODE-s2 resampling and differentiating, hereafter referred to as MLP proxy. And 3) in SPIN-ODE-s2, we propose to resample from the learned trajectories and then apply CDF. We evaluate the pre-training result both with and without dense interpolation. To ensure fairness, all methods without interpolation are trained for proportionally longer epochs to match the number of collocation points.
@@
TABLE II: Ablation studies. We report the quantitative results for our pre-training step on the AOXID dataset with or without key proposed components.
Table II presents the results on the AOXID dataset. The CDF method applied directly to sparse observations performs worst in both trajectory prediction and parameter estimation, reflecting its sensitivity to data resolution. While using the latent network output as a proxy for derivatives yields the best trajectory match with $\mathcal{L}_{y}=1.1\times 10^{-6}$, its higher $\mathcal{L}_{k}$ suggests that accurate reproduction of concentration trajectories does not necessarily imply reliable estimation of reaction coefficients. This discrepancy arises because neural ODE solvers perform integration over multiple internal steps between successive observation points, and the model receives no supervision at these intermediate time steps. Row 3 performs comparably to row 2, indicating no additional information from fitting the MLP without interpolation. Interpolation in trajectory space followed by differentiation can be interpreted as a smoothing operation in time-derivative space. The last row achieves the lowest rate coefficient error $\mathcal{L}_{k}=0.145$, indicating the most accurate recovery of system dynamics.
IV-D Robustness studies
In real-world chemistry experiments, measurement frequency is often limited by instrumentation constraints and experimental costs, resulting in sparse observations, for example, sometimes as few as 10 time points over an entire reaction. To evaluate the robustness of our method under real-world conditions, we simulate increased sparsity by downsampling the concentration trajectories of the AOXID problem by factors of 2, 4, and 10, while keeping the total time span fixed. We then assess the accuracy of the CRNN pre-training under these conditions.
As shown in Table III, our method remains stable and yields physically plausible estimates even as data sparsity increases. The trajectory loss $\mathcal{L}_{y}$ and coefficient error $\mathcal{L}_{k}$ increase gradually, but the degradation in performance remains moderate. Notably, even with a 10-fold reduction in observations, the model achieves a reasonable rate coefficient error of 0.45. These results suggest that our approach is robust to sparse supervision and may be applicable to low-resolution experimental datasets.
Downsampling rate ODE fitting Rate coeff. estimat

### v3

@@ 3.2 Proposed method
where $y_{\text{max}}$ and $y_{\text{min}}$ are the species-wise maximum and minimum concentrations used in the normalisation layer, and $t_{\text{scale}}$ is the time span of the entire measurements. The ODE solver integrates the output time derivative to a trajectory as Eq. (9), and the loss function is correspondingly normalised by the scale of concentrations as:
@@
In addition to the loss on predicted state variables, we introduce two derivative-based loss terms on the velocity and acceleration of concentration changes. These act as a stabilisation strategy to regularise the training under stiffness, rather than as strict physical constraints. Specifically, we compute the 1st-order and 2nd-order derivatives losses ($\mathcal{L}_{\dot{y}}$ and $\mathcal{L}_{\ddot{y}}$) with respect to time, then apply the same scaling technique as in Eq. (11), and the final loss function is:
$\mathcal{L}=\mathcal{L}_{y}+\alpha\mathcal{L}_{\dot{y}}+\beta\mathcal{L}_{\ddot{y}},$ (12)
where $\alpha$ and $\beta$ are both set to $0.1$.
SPIN-ODE step 2: Pre-training CRNN with estimated derivatives. While non-linear system dynamics can be approximated using simple feed-forward networks such as MLPs, the parameters learned by such black-box models typically lack interpretability. In contrast, physics-based neural ODE encodes parameters and mathematical operators derived directly from physical equations [22, 28]. Following [22, 43], we leverage the Chemical Reaction Neural Network (CRNN) (see Figure 2), a physics-based neural network tailored for discovering chemical reaction parameters. Given fixed reaction pathways, CRNN reformulates the rate law in Eq. (2) as a linear combination of the logarithm of learnable rate coefficients $\theta_{i}$ and the logarithms of species concentrations, followed by an exponential transformation:
$r_{i}=\exp\left(\ln(\theta_{i})+s_{i,a}^{f}\ln[A]+s_{i,b}^{f}\ln[B]+s_{i,c}^{f}\ln[C]\right).$ (13)
Each CRNN node corresponds to an individual reaction, and nodes can be stacked to represent a reaction system. The overall time derivative of species concentrations $d\mathbf{Y}/dt$, is obtained by linearly transforming the reaction rate vector $\mathbf{R}$, through the net stoichiometric matrix $\mathbf{S}$, defined as $s_{i,j}=-s_{i,j}^{f}+s_{i,j}^{r}$:
@@
To facilitate CRNN training, we deviate from prior work in which the CRNN output is directly integrated and trained end-to-end on concentration trajectories. Such coupling is prone to instability in stiff regimes, since small integration errors can rapidly accumulate during back-propagation. Instead, in SPIN-ODE we decouple the CRNN from ODE integration and train it as a supervised model of the differential equations as in Eq. (8). Specifically, we first generate paired data of concentration and time derivative from the fitted trajectories in step 1: when the black-box neural ODE fits the concentration trajectory ${{y}(t_{i})}_{i=1}^{M}$, we increase the temporal resolution by interpolation to obtain ${\hat{y}(t_{i})}_{i=1}^{N}$ with $N>M$. We then approximate the derivatives $\dot{\hat{y}}(t)$ at state $\hat{y}(t)$ using finite difference:
$\dot{\hat{y}}(t_{i})\approx\begin{cases}\dfrac{\hat{y}(t_{i+1})-\hat{y}(t_{i-1})}{t_{i+1}-t_{i-1}},&1<i<N,\\ \dfrac{\hat{y}(t_{2})-\hat{y}(t_{1})}{t_{2}-t_{1}},&i=1,\\ \dfrac{\hat{y}(t_{N})-\hat{y}(t_{N-1})}{t_{N}-t_{N-1}},&i=N.\end{cases}$ (15)
@@
SPIN-ODE-s1 $2.8\times 10^{-3}$ - $1.1\times 10^{-5}$ - $1.4\times 10^{-5}$ -
@@
Table 1: Overall comparisons between our proposed three-step SPIN-ODE framework and Ji et al. [22] model. We show the scaled MSE loss of the ODE trajectory fitting, as well as the mean absolute error on log-scaled rate coefficient estimation. - means the network involved in the current step does not provide estimation, while * means runtime error.
This design offers two main benefits. First, by training CRNN directly on interpolated and thus smoothed (y^,y^˙)($\hat$$\dot{\hat{y}}$) pairs rather than coupling it with an ODE solver, we avoid unstable integration during the early stage of CRNN training. Once the parameters are reasonably initialised, ODE integration can be safely reintroduced in step 3 for fine-tuning. Second, keeping the CRNN parameters aligned with physical rate laws preserves interpretability, which would be compromised if the scaling techniques from step 1 were applied directly. In that case, reaction rates would be non-linearly distorted, and the learned coefficients would lose their physical meaning.
SPIN-ODE step 3: Fine-tuning CRNN with ODE integration. After pre-training, the CRNN is coupled with a stiff ODE solver (Eq. (9)) to further refine the rate coefficients. Unlike step 2, where supervision is limited to discrete $(\hat{y}(t_{i}),\dot{\hat{y}}(t_{i}))$ pairs, the solver integrates through intermediate time steps, thereby constraining the dynamics between sampled points and enforcing consistency across the full trajectory. Eventually, the optimised parameters $\theta_{i}$ from CRNN are taken as the estimated reaction rate coefficients $k_{i}$.
4 Experiments
@@
4.1 Experimental setups
Our method is first validated using the Robertson problem, a classical benchmark comprising three chemical species and three stiff reactions as presented in Eq. (1). The governing equations for the Robertson problem are explicitly defined in Eq. (5), with ground-truth reaction rate coefficients as $0.04,3\times 10^{7},1\times 10^{4}]$.
Real-world atmospheric chemistry systems typically involve significantly more complex dynamics, exhibiting substantial stiffness across numerous reactions. Consequently, we further evaluate our method on the POLLU problem [45], a widely utilised atmospheric chemistry model for stiff ODE benchmarks. This system includes 25 reactions among 20 chemical species, with reaction rate coefficients spanning from $10^{-3}$ to $10^{12}$. In addition, we include an autoxidation reaction (AOXID) scheme involving 44 species and 49 reactions. It is a simplified system representing realistic autoxidation processes in the atmosphere characterised by intricate radical chain mechanisms [38]. The reaction rate coefficients in this system vary dramatically, ranging from approximately $10^{-14}$ to $10^{6}$. Comprehensive details for the POLLU and AOXID reaction schemes are provided in the supplementary materials [37].
Datasets for the Robertson and POLLU problems are numerically generated using a stiff-aware Backward Differentiation Formula (BDF) solver available in the SciPy library. The AOXID dataset is generated using the Rosenbrock solver within the well-known Kinetic PreProcessor (KPP) [9]. For the Robertson system, initial concentrations are set as $y_{1},y_{2},y_{3}]=[1,0,0]$, with 50 data points sampled logarithmically spaced between $10^{-5}$ and $10^{5}$ seconds. For the POLLU and AOXID datasets, initial conditions detailed in the supplementary material are used, and 100 data points are uniformly sampled across the time span of 0.1 and 100 seconds, respectively. To mitigate numerical instabilities arising from prolonged integration and accumulation of errors, a sliding window approach of size 20 and stride of 10 is adopted during the training for POLLU and AOXID.
The proposed SPIN-ODE leverages the stiff-aware differentiable ODE solver Kvaerno3 from the diffrax library [24]. For the POLLU dataset, the MLP used in step 1 has an input and output dimension of 20 and a hidden layer size of 128, resulting in a total of 21,780 trainable parameters. In contrast, the CRNN in steps 2 and 3 has only 25 trainable parameters, corresponding to the number of reactions in the system. Network parameters are optimised using the Adam algorithm, initialised with a learning rate of 0.001, which is annealed when training loss does not decrease for 10% of the total epochs. All experiments were conducted on a single GPU node of the CSC Mahti supercomputer, using one NVIDIA A100-SXM4-40GB GPU and 16 cores of an AMD Rome 7H12 CPU. For the POLLU dataset, the runtimes for the three stages of SPIN-ODE were 1.17 h, 6.26 min, and 11.83 min, respectively. Convergence curves for all three training stages are shown in Figure 1 of the supplementary material [37]. The code is available at https://github.com/pvvq/SPIN-ODE, with an archived version in [36].
4.2 Comparison with state of the art
Our overall comparisons include both ODE trajectory fitting and rate coefficient estimation. Given the wide-ranging magnitudes of the chemical concentration and reaction rate coefficients, we evaluate the model estimation errors on a normalised scale. To the best of our knowledge, this is the first study to address stiff chemical ODEs for rate coefficient estimation. The only closely related work is by Ji et al. [22], which we reimplemented for comparison.
The quantitative results on ODE fitting and rate coefficients are reported in Table 1. We show our results in every step (SPIN-ODE-s1, SPIN-ODE-s2, SPIN-ODE-s3). As shown in Table 1, ours outperforms Ji et al. [22] in every step. We note that Ji et al. fail to converge on the AOXID dataset, and a large error on the POLLU dataset. On the other hand, our final results (SPIN-ODE-s3) achieve low errors in rate coefficient estimation and ODE fitting, specifically $7.4\times 10^{-3}$ for Robertson, $8.6\times 10^{-6}$ for POLLU, and $4.5\times 10^{-6}$ for AOXID. The promising results on the AOXID dataset indicate that our approach could be used in real-world atmospheric chemistry. Furthermore, by comparing SPIN-ODE-s3 with SPIN-ODE-s2, it demonstrates the benefit from the final refinement process. Note that in SPIN-ODE-s1, we use MLPs to implicitly learn ODE fitting and it cannot estimate rate coefficients yet. Compared to SPIN-ODE-s1, SPIN-ODE-s2 exhibits larger ODE fitting errors. These errors result from the oscillatory nature of derivatives in stiff systems, and training on smoothed derivatives captures only the slower components rather than the full system dynamics. This further indicates the necessity of the third fine-tuning step.
To better visualise the ODE trajectory fitting, we plot the trajectory predictions in Figure 3. From left to right, we have the results for Robertson, POLLU and AOXID. Note that we select the three most representative reactions from POLLU and AOXID for visualisation. The complete reaction plots can be found in the supplementary material. Since the method by Ji et al. performs reliably only on the Robertson problem, we did not show its trajectories on POLLY and AOXID. For our approach, it fits the ground truth trajectories in all three cases.
Since our primary focus is on rate coefficient estimation, we also plot the scatter diagram to demonstrate all rate coefficients in Figure 4. We can see that the model predictions (blue cross) closely match with ground truth values (yellow circles). Note that for the POLLU problem, coefficients 15-19 correspond to hydroxyl (OH) cycling reactions, whose rate coefficients have been intensively studied and are provided during training to facilitate convergence, and thus excluded from the estimation result.
4.3 Ablation studies
@@
We compare three alternative strategies against our proposed SPIN-ODE-s2 with interpolation to highlight its design choices: (1) applying finite differences directly to the original observations to obtain time derivatives, without SPIN-ODE-s1 MLP fitting (referred to as Direct FD); (2) using the MLP outputs from SPIN-ODE-s1 to train the CRNN, without resampling or differentiating (referred to as MLP proxy); and (3) applying finite differences to resampled trajectories as in SPIN-ODE-s2, but without interpolation. To ensure fairness, all methods without interpolation are trained for proportionally more epochs to match the number of collocation points.
@@
Table 2: Ablation studies. We report the quantitative results for our pre-training step on the AOXID dataset with or without key proposed components.
Table 2 presents the results on the AOXID dataset. Applying finite differences directly to sparse observations performs worst in both trajectory prediction and parameter estimation, reflecting strong sensitivity to data resolution. Using the black-box MLP output as a proxy for derivatives yields the best trajectory match ($\mathcal{L}_{y}=1.1\times 10^{-6}$), but its higher rate coefficient error shows that reproducing trajectories alone does not guarantee reliable parameter estimation, consistent with our earlier discussion of the disparity between system dynamics revealed by concentrations and their derivatives. The variant without interpolation performs similarly to the MLP proxy, indicating that resampling alone provides little benefit. By contrast, incorporating interpolation before finite differencing yields the lowest rate coefficient error ($\mathcal{L}_{k}=0.145$), making it the most effective training strategy among the tested variants.
4.4 Robustness studies
In real-world chemistry experiments, measurement frequency is often constrained by instrumentation and cost, leading to sparse observations, sometimes with as few as 10 time points for an entire reaction. To evaluate the robustness of our method under such conditions, we simulate data sparsity by downsampling the AOXID concentration trajectories by factors of 2, 4, and 10, while keeping the total time span fixed. We then assess the accuracy of CRNN pre-training under these scenarios.
As shown in Table 3, our method remains stable and produces physically plausible estimates even as data sparsity increases. The trajectory loss $\mathcal{L}_{y}$ and coefficient error increase gradually, but the degradation in performance remains moderate. Even with a 10-fold reduction in observations, the model achieves a reasonable rate coefficient error of 0.45, demonstrating robustness to sparse supervision and applicability to low-resolution experimental datasets.
Downsampling rate ODE fitting Rate coeff. estimate